what would a 5.521 helix look like compared to a pi helix?
Answer #1

Generally, pi-helix will be embedded in a longer alpha-helix with some more hump compared the alpha-helix.

Now, 5.521/pi = 5.521/(22/7) = 1.757

Therefore, 5.521 helix = 1.75668 * pi-helix

Therefore, 5.521-helix will also be embedded in a longer alpha-helix with more hump compared to pi-helix as well as alpha-helix.